How much pure alcohol must be added to 50 gallons of gasohol that is 10% alcohol to make gasohol that is 20% alcohol?
Answer by mananth(16946)

How much pure alcohol must be added to 50 gallons of gasohol that is 10% alcohol to make gasohol that is 20% alcohol?
...
pure alcohol ------------100%---------x gallons
solution ----------------10%----------50 gallons of gasohol
gasohol---------------- 20%-----------x+50 gallons
..
x*1+50*0.1=(x+50)*0.2
multiply by 10
10x+50=2(x+50)
10x+50=2x+100
8x=50
/8
x= 50/8
x= 6.25 gallons of pure alcohol.
...
m.ananth@hotmail.ca